Explain what happens to the potential difference across a discharging capacitor after three (3) time constants have elapsed. (3 marks)
1
Expert's answer
2020-04-27T10:14:20-0400

Formula for potential difference as a function of time if given by


"V_c=V_s\times e^{\frac{-t}{RC}}"

  • "V_c"  is the voltage across the capacitor
  • "V_s" is the supply voltage
  • "t"  is the elapsed time since the removal of the supply voltage
  • "RC"  is the time constant of the RC discharging circuit

So here t= 3RC

putting the value

We get,


"V_c= V_s\times e^{-3}"

So after this time it will reduced "\frac{1}{e^3}" of its original value.